OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2x - 2 <= 4? 3 x x 3 A number line with a closed circle on 3 and shading to the right. A number line with a closed circle on 3 and shading to the left.

OpenStudy (anonymous):

2x - 2 <= 4 2x <= 4 + 2 2x <= 6 2x/2 <= 6/2 x <= 3 This question does not make sense.....none of these answer choices represent the solution.

OpenStudy (anonymous):

oops....made a mistake....the last one REPRESENTS the solution.